An isosceles triangle in which angles B and C have the same measure. Find the measure of each angle whose degree is represented with variables.
A= 4x + 2y
B= 9y - 2
C= 10x + 10
The assignment calls for setting it up as a system of equations and solving the problem. I have no idea where to even begin.
I tried something like 4x + 2y = 180- 9y -2 + 10x + 10
that definately was not right! Lol!
I would appreciate any help.

Answer by Greenfinch(383)

We know B = C,so we know 10x + 10 = 9y - 2
thus x = (9y -12)/10
and A + B + C = 180
14x + 11y + 8 = 180
14 (9y - 12)/10 + 11y = 172
126y - 168 + 110y = 1720
236y = 1888
y = 8
x = (72 - 12)/10 = 6
So B = 70
C = 70
A = 40